1. What is the mass in grams of a single atom of Thorium? 
  2. What mass of Thallium has the same number of atoms as 65.2 g of Potassium?
1
Expert's answer
2022-05-23T17:02:05-0400

1.

a) According to Mendeleev's periodic table of chemical elements, the relative atomic mass of thorium (Th) is 232,4 gr . this value refers to 1 mole of thorium (Th).

b) According to Avagadro's law, 1 mole of any substance contains 6,02 * 1023 particles (atoms, molecules). so to find a single atomic mass, we divide the relative atomic mass by 6,02 * 1023 :

232,4 / 6,02 * 1023 = 38,6 * 10-23

ANSWER: mass of a single atom of Thorium is 38,6 * 10-23 gr or 3,86 * 10-24 gr.

2.

Mr (Tl) = 204,4 gr

Mr (K) = 39 gr

a) We find the moles of 65.2 g of potassium. for this we divide 65.2 by the relative atomic mass of potassium:

n = 65,2 / 39 = 1,67 mol

According to Avagadro's law, the number of particles (atoms, molecules) in one mole of any substance is the same. so the number of atoms in 1,67 mol potassium (K) is equal to the number of Thallium (Th) atoms in the same mole. hence ,the amount of thallium is 1.67 mol. we find thallium mass. for this we divide 1.67 by the relative atomic mass of thallium: 1,67 * 204,4 = 341,35 gr.

ANSWER: 341,35 gr.